40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14397     Entry time: Fri Jan 11 16:38:57 2019
Author: gautam 
Type: Update 
Category: General 
Subject: Some alignment checks 

The pumpdown seems to be progressing smoothly, so I think we are going to stick with the plan decided on Wednesday, and vent the IFO on Monday at 8am. I decided to do some checks of the IFO alignment.

I turned on the PSL again (so goggles are advisable again inside the VEA until this work is done), re-locked the PMC, and opened the PSL shutter into the vacuum (still low power 100 mW beam going into vacuum). The IMC alignment required minor tweaking, but I recovered ~1300 cts transmission which is what it was --> so we didn't macroscopically change the input pointing into the IMC while working on the IOO table.

Centering the ITMY oplev spot, there is a spot on the AS camera roughly centered on the control room monitor, so the TT pointing must also be pretty close.

Then I centered the ITMY oplev spot to check how well-aligned or otherwise the Michelson was - the BS has no Oplev so there was considerable angular motion of the Michelson spot, but it looked like on average, it was swinging around through a well aligned place. I saved the slow bias voltages for the ITMs and BS in this config.

Then I re-aligned ETMX and checked the green transmission - it was okay, ~0.3, and I was able to increase it to ~0.4 using the EX green PZT mirrors. So far so good.

Finally, I tried to lock the X-arm on IR - after zeroing the offsets on the transmission QPD, there seems to be a few flashes as the cavity swings through resonances, but no discernible PDH error signal. Moreover the input pointing of the IR into the X arm is controlled by the BS which is swinging around all over the place right now, so perhaps locking is hopeless, but the overall alignment of the IFO seems not too bad. Once ETMY is cleaned and put back in place, perhaps the Y arm can be locked.

I shuttered the PSL and inserted a manual beam block, and also turned off the EX laser so that we can vent on Monday without laser goggles.

*Not directly related to this work: we still have to implement the vacuum interlock condition that closes the PSL shutter in the event of a vacuum failure. It's probably fine now while the PSL power is attenuated, but once we have the high power beam going in, it'd be a good to revert to the old standard.

Attachment 1: pd82.png  151 kB  | Hide | Hide all
pd82.png
Attachment 2: LSC_X.png  57 kB  Uploaded Fri Jan 11 18:15:31 2019  | Hide | Hide all
LSC_X.png
ELOG V3.1.3-